For each of the following numbers, find the smallest whole number by which it should be multiplied so as to get a perfect square number. Also find the square root of the square number so obtained.  ii. 180

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student
180 = 2×2×3×3×5
= (2×2)×(3×3)×5
Here, 5 cannot be paired.
∴ We will multiply 180 by 5 to get perfect square.Newnumber=180×5=900

900 = 2×2×3×3×5×5×1
⇒ 900 = (2×2)×(3×3)×(5×5) ⇒ 900 = 22×32×52
⇒ 900 = (2×3×5)2⇒√900=2×3×5=30
Thanks